OpenStudy (anonymous):

algebra 2 help please!!! Solve for x: 4 over x plus 4 over quantity x squared minus 9 equals 3 over quantity x minus 3. Solve for x: 2 over quantity x plus 2 plus 1 over 5 equals 6 over quantity x plus 5. Classify the solutions of 1 over x plus 4, plus one half, equals 1 over x plus 4 as extraneous or non-extraneous. Solve for x: 3 over 3 x plus 1 over quantity x plus 4 equals 10 over 7x .

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@Leader

zepdrix (zepdrix):

\[\Large\rm \frac{4}{x}+\frac{4}{x^2-9}=\frac{3}{x-3}\]This is our problem to solve?

OpenStudy (anonymous):

i should have updated that one i got that answer

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Solve for x: 2 over quantity x plus 2 plus 1 over 5 equals 6 over quantity x plus 5. \[\Large\rm \frac{2}{(x+2)}+\frac{1}{5}=\frac{6}{(x+5)}\]Mmmm ok something like this?

zepdrix (zepdrix):

We would like to multiply both sides by the `Least Common Multiple` of the denominators. It's pretty much the same process as finding a common denominator, but we're multiplying through by that denominator so it cleans things up.

OpenStudy (anonymous):

ok and yea thats the right one

zepdrix (zepdrix):

Our denominators have nothing in common, so our LCM is going to be \(\Large\rm (x+2)\cdot5\cdot(x+5)\)

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

\[\rm 5(x+2)(x+5)\left(\frac{2}{(x+2)}+\frac{1}{5}\right)=\left(\frac{6}{(x+5)}\right)5(x+2)(x+5)\]We'll multiply both sides by these denominators. It will get rid of all of the fractions for us.

zepdrix (zepdrix):

So what do we get for our first term? \[\Large\rm 5(x+2)(x+5)\cdot \frac{2}{x+2}=?\] Can you figure it out? :)

OpenStudy (anonymous):

im looking for the LCD? cause i got x+2

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Ah sorry I ran off :c busy morning. Least Common Multiple? The first fraction has (x+2), the second one has a 5, the last one has an (x+5). So our LCM is the product of all three of those, since they have nothing in common.

zepdrix (zepdrix):

\[\rm 5(x+2)(x+5)\left(\frac{2}{(x+2)}+\frac{1}{5}\right)=\left(\frac{6}{(x+5)}\right)5(x+2)(x+5)\]So I'm multiplying both sides by the LCM. Distributing it to each fraction gives us:\[\rm \frac{2\cdot5(x+2)(x+5)}{(x+2)}+\frac{5(x+2)(x+5)}{5}=\frac{6\cdot5(x+2)(x+5)}{(x+5)}\]

zepdrix (zepdrix):

From there, we have some nice cancellations\[\rm \frac{2\cdot5\cancel{(x+2)}(x+5)}{\cancel{(x+2)}}+\frac{\cancel{5}(x+2)(x+5)}{\cancel{5}}=\frac{6\cdot5(x+2)\cancel{(x+5)}}{\cancel{(x+5)}}\]

zepdrix (zepdrix):

Which leaves you with:\[\Large\rm 10(x+5)+(x+2)(x+5)=30(x+2)\]No more fractions, yay!

OpenStudy (anonymous):

its all good and what do i do from here?

zepdrix (zepdrix):

Multiply out the brackets, get everything on the same side, combine like-terms.

zepdrix (zepdrix):

Boy these are longs problems :O flvs so mean hehe

OpenStudy (anonymous):

hence my name lol

OpenStudy (anonymous):

im not good at multiplying the like terms especially when there are so many, my dyslexia kicks in can you help out?

zepdrix (zepdrix):

Distribute the 10, what do you get for the first set of terms?

OpenStudy (anonymous):

like distribute to the 30?

OpenStudy (anonymous):

or among the brackets?

zepdrix (zepdrix):

\[\Large\rm \color{orangered}{10(x+5)}+(x+2)(x+5)=30(x+2)\]Just this orange part to start. What two terms do you get out of that when you give the 10 to the x and 5?

OpenStudy (anonymous):

oh, i just did the whole thing and got 0 and 13

zepdrix (zepdrix):

0 and 13? Yayyyy good job \c:/

OpenStudy (anonymous):

SWEET!!!!!

OpenStudy (anonymous):

can you help me with the next one?

zepdrix (zepdrix):

Sorry gotta get ready for work :c